OpenStudy (anonymous):

Evaluate log base 9 of 27 Can someone help explain this to me?

OpenStudy (zzr0ck3r):

\(\log_9(27)=\log_{3^2}(27)\) So \((3^2)^x=27\) i.e. \(2*x=3\implies x=\dfrac{3}{2}\)

OpenStudy (anonymous):

Is that the answer? Sorry I can't find this type of question in my notes :(

OpenStudy (zzr0ck3r):

Which part of my explanation do you not understand?

OpenStudy (anonymous):

Why do you square 3 and where does x come from?